A man is four times as ols as his son.in five years he will be three times as old. Find their present ages
Answer by algebrahouse.com(1659)

x = son's age now
4x = man's age now

x + 5 = son's age in 5 years
4x + 5 = man's age in 5 years

4x + 5 = 3(x + 5) {in 5 years, the man will be 3 times the son}
4x + 5 = 3x + 15 {used distributive property}
x + 5 = 15 {subtracted 3x from each side}
x = 10 {subtracted 5 from each side}
4x = 40 {substituted 5, in for x, into 4x}

son is 10
man is 40

To ask a question, visit: www.algebrahouse.com